solve each equation by completing the square p^2-5p= -4

Answer by ewatrrr(24785)


Hi
solving by completing the square
p^2-5p = -4
(p-5/2)^2 = -4 + 25/4
(p-5/2)^2 = 9/4
(p-5/2) = ± 3/2
p = 5/2 ± 3/2
p = 4 or p = 1
CHECKING our Answer***
p^2-5p +4 = 0
(p-4)(p-1) = 0
(p-4)= 0 p = 4
(p-1)=0 p = 1